OpenStudy (anonymous):

Finish the following table for the given function. x g(x) 3 5 7 9 g(x)=x^2 Start the independent variable x at 3 and use an increment of 2. a. 6, 1, 1, 9 b. 1, 1, 9, 25 c. 6, 10, 14, 18 d. 9, 25, 49, 81

OpenStudy (misty1212):

\[3^2=?\\ 5^2=?\\ 7^2=?\\ 9^2=?\]